OpenStudy (anonymous):

A life insurance company classifies its customers as being either high risk or low risk. If 20% of the customers are high risk, and high risk customers are three times more likely to file a claim, hat percentage of claims that are filed come from high risk customers?

OpenStudy (anonymous):

H + L = 100% customers H = 20%, L = 80% 3 H + L = claims (3)(20) + (1)(80) = claims 60 + 80 = 140 60/140 due to H and 80/140 due to L make fractions into % for answer.

OpenStudy (anonymous):

Thanks - I knew this was an easy problem, but I just wasn't sure where to start.

OpenStudy (anonymous):

Okay, I have another one here: Suppose that \[X _{1},...,X _{100} \] are random variables with \[E(X _{i})=100, E(X _{i}^{2})=10100\] If \[Cov(X _{i},X _{j})=-1, i \neq j\] what is Var(S), where\[S=\sum_{i=1}^{100}X _{i}\]? I got Var(S) to be 100, which is the correct answer but what good does knowing the covariance do? I tried using it, and it just throws everything off.